OpenStudy (anonymous):

Calculate the length of the catenary x from 0 to b

OpenStudy (anonymous):

Sir give me advice :(

OpenStudy (anonymous):

wht is function here

OpenStudy (anonymous):

f(x) = a cosh (x/a) with vertex (0,a).

OpenStudy (anonymous):

ds=sqrt(dx^2+dy^2) y=acosh(x/a) dy=sinh(x/a)dx so s= [integral from 0 to b] sqrt(1+sinh(x/a)^2)dx s= [integral from 0 to b] cosh(x/a)dx so =sinh(b)

OpenStudy (anonymous):

a sinh b/a

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sir, Is a catenary a type of canary ? gummy bears :) leave the fig trees alone and come back to openstudy

OpenStudy (lalaly):

come baaaaaaaaaackkkkkkkkkkkkkkk

OpenStudy (anonymous):

Hola estudier, come back, please ! everyone misses you. :-)

OpenStudy (lalaly):

i hope he gets them